# On new identities of Jacobi sums and related cyclotomic matrices

### v1

@@ Abstract.
In this paper, using some arithmetic properties of Jacobi sums, we investigate some products involving Jacobi sums and reveal the connections between these products and certain cyclotomic matrices. In particular, as an application of our main results, we confirm a conjecture posed by Z.-W. Sun in 2019.
@@ 1.1. Notation
$\phi(x)=\begin{cases}1&\mbox{if}\ x\in\mathcal{S}_{q}\setminus\{0\},\\ 0&\mbox{if}\ x=0,\\ -1&\mbox{if}\ x\in\mathbb{F}_{q}\setminus\mathcal{S}_{q}\end{cases}$
@@ 1.2. Background and motivation
be the projective Fermat curve defined over $\mathbb{F}_{q}$, where $\mathbb{F}_{q}^{{\rm alg}}$ is an algebraic closure of $\mathbb{F}_{q}$, and $\mathbb{P}^{2}(\mathbb{F}_{q})$ is the projective plane. Consider the zeta function of $C_{n}$ defined by
@@
$\zeta_{C_{m}}=\zeta_{C_{m}}(t)=\frac{P_{m}(t)}{(1-t)(1-qt)}.$
Weil proved that the integral polynomial $P_{m}(t)$ has close relations with products of Jacobi sums, that is,
@@ 1.3. Main results
As an application of Theorem 1.2, we can confirm this conjecture.
@@
Conjecture 1.1 holds.
@@ 2.1. On a variant of the Gauss lemma
Now we are in a position to prove the following variant of Gauss lemma, which will play an important role in the proof of main theorems.
@@ 2.4. Some necessary lemmas
(ii) Moreover, if $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, then for $k=1,2,\cdots,n$, we have
@@
$\chi_{q}^{k}(d)\cdot\lambda_{k}\in\mathbb{R}{\bm{i}}:=\left\{x{\bm{i}}:\ x\in\mathbb{R}\right\},$
@@
Let $q=2n+1\equiv 1\ ({\rm{mod}}\ 4)$ be an odd prime power. Then for any $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, the $n+1$ numbers $\pm\sqrt{-n},\lambda_{1}(d),\lambda_{2}(d),\cdots,\lambda_{n-1}(d)$ are exactly all the eigenvalues of $T_{q}(d)$.
@@
3. Proofs of Theorem 1.1
Proof of Theorem 1.1. Let $q\equiv 3\ ({\rm{mod}}\ 4)$ be an odd prime power with n=(q1)/2n=(q-1)/2. We first prove that
@@
$S_{j}(a):=\left\{k\in(0,n/2):\ \frac{j-1}{2}n<\{ka\}_{2n}<\frac{j}{2}n\right\}.$
@@
We next evaluate $x_{q}^{2}$. Fix a generator $g$ of cyclic group $\mathbb{F}_{q}^{\times}$ and let $s_{j}=g^{2j}$ for $j=0,1,\cdots,n-1$. Then the matrix
@@
4. Proofs of Theorem 1.2 and its corollary
@@
$S_{j}(a):=\left\{k\in(0,n/2):\ \frac{j-1}{2}n<\{ka\}_{2n}<\frac{j}{2}n\right\}.$
@@
Now we turn to the proof of our last result.
@@
is a square of the integer $z_{p}$. Moreover, Z.-W. Sun [11, Theorem 1.2(iii)] proved that
$\left(\frac{\det T_{d}(p)}{p}\right)=1$
whenever $d\in\mathbb{F}_{p}^{\times}\setminus\mathcal{S}_{p}$. This implies that $\det T_{d}(p)\in\mathbb{Z}\setminus\{0\}$.
In view of the above, $|\det T_{p}(d)|/(2^{(p-1)/2}b_{p})$ is indeed a positive integral square not depending on $d$. This completes the proof. ∎
@@ 5. Concluding Remarks
for some intger $t_{p}$ with $(\frac{t_{p}}{p})=1$. Moreover, by the proof of Corollary 1.1, we have proven that
@@
$z_{p}=\frac{{\bm{i}}^{\frac{p-5}{4}}\cdot I_{p}(\chi_{p})\cdot\sqrt{p-1}}{2^{(p-5)/4}}\in\mathbb{Z}.$
In view of the above, one can verify that Conjecture 5.1 holds if and only if the rational number $4z_{p}/(p-1)$ is not only an integer but also a quadratic residue modulo $p$. However, we can not prove this result by the methods used in this paper.

### v2

@@ Abstract.
In this paper, using some arithmetic properties of Jacobi sums, we investigate some products involving Jacobi sums and reveal the connections between these products and certain cyclotomic matrices. In particular, as an application of our main results, we confirm a conjecture posed by Z.-W. Sun in 2019, and obtain a stronger result.
@@ 1.1. Notation
$\phi(x)=\begin{cases}1&\mbox{if}\ x\in\mathcal{S}_{q}\setminus\{0\},\\ 0&\mbox{if}\ x=0,\\ -1&\mbox{if}\ x\in\mathbb{F}_{q}\setminus\mathcal{S}_{q}.\end{cases}$
@@ 1.2. Background and motivation
be the projective Fermat curve defined over $\mathbb{F}_{q}$, where $\mathbb{F}_{q}^{{\rm alg}}$ is an algebraic closure of $\mathbb{F}_{q}$, and $\mathbb{P}^{2}(\mathbb{F}_{q}^{{\rm alg}})$ is the projective plane. Consider the zeta function of $C_{n}$ defined by
@@
$\zeta_{C_{m}}(t)=\frac{P_{m}(t)}{(1-t)(1-qt)}.$
Weil proved that the integral polynomial $P_{m}(t)$ has close relations with certain product of Jacobi sums, that is,
@@ 1.3. Main results
As an application of Theorem 1.2, we can confirm this conjecture and obtain a stronger result.
@@
Let $p\equiv 1\ ({\rm{mod}}\ 4)$ be a prime and let notations be as above. Set
$z_{p}:=\frac{y_{p}}{2^{(n-2)/2}}.$
Then $z_{p}$ is an integer such that
$\left(\frac{z_{p}}{p}\right)=1,$
and
$|\det T_{p}(d)|/(2^{(p-1)/2}b_{p})=z_{p}^{2}.$
@@ 2.1. On a variant of the Gauss lemma
Now we are in a position to prove the following variant of Gauss lemma, which will play an important role in the proof of our second theorem.
@@ 2.4. Some necessary lemmas
(ii) Moreover, if $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, then for any $k=1,2,\cdots,n$, we have
@@
$\chi_{q}^{k}(d)\cdot\lambda_{k}\in\mathbb{R}{\bm{i}}=\left\{x{\bm{i}}:\ x\in\mathbb{R}\right\},$
@@
Let $q=2n+1$ be an odd prime power with $q\equiv 1\ ({\rm{mod}}\ 4)$. Then for any $d\in\mathbb{F}_{q}^{\times}\setminus\mathcal{S}_{q}$, the $n+1$ numbers $\pm\sqrt{-n},\lambda_{1}(d),\lambda_{2}(d),\cdots,\lambda_{n-1}(d)$ are exactly all the eigenvalues of $T_{q}(d)$.
@@
3. Proofs of Theorems 1.11.2
Proofs of Theorems 1.1. Let $q\equiv 3\ ({\rm{mod}}\ 4)$ be an odd prime power with n=(q1)/2n=(q-1)/2. We first prove that
@@
$S_{j}(a)=\left\{k\in(0,n/2):\ \frac{j-1}{2}n<\{ka\}_{2n}<\frac{j}{2}n\right\}.$
@@
We next evaluate $x_{q}^{2}$. Fix a generator $g$ of cyclic group $\mathbb{F}_{q}^{\times}$ and let $s_{j+1}=g^{2j}$ for $j=0,1,\cdots,n-1$. Then the matrix
@@
We next turn to the proof of our second theorem.
@@
$S_{j}(a)=\left\{k\in(0,n/2):\ \frac{j-1}{2}n<\{ka\}_{2n}<\frac{j}{2}n\right\}.$
@@
4. Proof of Corollary 1.1
Now we turn to the proof of our last result. We first introduce the following lemma.
Lemma 4.1.
Let $p\equiv 1\ ({\rm{mod}}\ 4)$ be a prime and let
$B_{p}=\prod_{k=1}^{(p-1)/4}\binom{2k}{k}$.
Then
$\left(\frac{B_{p}}{p}\right)=\left(\frac{2}{p}\right).$
Proof.
By the definition of binomial coefficients, one can verify that
(4.1) $\left(\frac{B_{p}}{p}\right)=\prod_{k=1}^{(p-1)/4}\left(\frac{(2k)!}{p}\right)=\prod_{x=1}^{(p-1)/2}\left(\frac{x}{p}\right)^{\frac{p-1}{4}}\left(\frac{x}{p}\right)^{\lfloor\frac{x-1}{2}\rfloor}=(-1)^{\frac{p-1}{4}}\prod_{x=1}^{(p-1)/2}\left(\frac{x}{p}\right)^{\lfloor\frac{x-1}{2}\rfloor}.$
For $i=0,1,2,3$, let
$N_{i}(p)=\#\left\{x\in(0,p/2):\ \left(\frac{x}{p}\right)=-1\ \text{and}\ x\equiv i\ ({\rm{mod}}\ 4)\right\}.$
As $p\equiv 1\ ({\rm{mod}}\ 4)$, one can verify that
$N_{0}(p)+N_{1}(p)$
$=$ $\#\left\{x\in(0,p/8):\ \left(\frac{x}{p}\right)=-1\right\}+\#\left\{x\in(p/2,p):\ \left(\frac{x}{p}\right)=-1\ \text{and}\ x\equiv 0\ ({\rm{mod}}\ 4)\right\}$
$=$ $\#\left\{x\in(0,p/8):\ \left(\frac{x}{p}\right)=-1\right\}+\#\left\{x\in(p/8,p/4):\ \left(\frac{x}{p}\right)=-1\right\}$
$=$ $\#\left\{x\in(0,p/4):\ \left(\frac{x}{p}\right)=-1\right\},$
and that
$N_{0}(p)+N_{2}(p)=\#\left\{x\in(0,p/4):\left(\frac{x}{p}\right)=-\left(\frac{2}{p}\right)\right\}.$
This implies
(4.2) $N_{1}(p)+N_{2}(p)\equiv\frac{p-1}{4}\ ({\rm{mod}}\ 2).$
Using $p\equiv 1\ ({\rm{mod}}\ 4)$ again, we obtain
(4.3) $N_{0}(p)+N_{1}(p)+N_{2}(p)+N_{3}(p)=\frac{p-1}{4}.$
By (4.2) and (4.3), we see that
$N_{0}(p)+N_{3}(p)\equiv 0\ ({\rm{mod}}\ 2).$
By this and (4.1), we obtain
$\left(\frac{B_{p}}{p}\right)=(-1)^{\frac{p-1}{4}+N_{0}(p)+N_{3}(p)}=(-1)^{\frac{p-1}{4}}=\left(\frac{2}{p}\right).$
This completes the proof. ∎
To state our next lemma, we need introduce the following notations. Let $\zeta_{p-1}$ be a primitive (p1)(p-1)-th root of unity. Let the cyclotomic field $K=\mathbb{Q}(\zeta_{p-1})$ and let $\mathcal{O}_{K}$ be the ring of all algebraic integers over $K$. Let $\mathfrak{p}$ be a prime ideal of $\mathcal{O}_{K}$ with $p\in\mathfrak{p}$. Then the Teichümuller character $\omega_{\mathfrak{p}}$ of $\mathfrak{p}$ is defined by
$\omega_{\mathfrak{p}}\left(x\ {\rm mod}\ \mathfrak{p}\right)\equiv x\ ({\rm{mod}}\ \mathfrak{p})$
for any $x\in\mathcal{O}_{K}$. As $p$ splits totally over $\mathcal{O}_{K}$, we have
$\mathcal{O}_{K}/\mathfrak{p}\cong\mathbb{F}_{p}.$
Thus, if we identify $\mathcal{O}_{K}/\mathfrak{p}$ with $\mathbb{F}_{p}$, then it is easy to verify that $\omega_{\mathfrak{p}}$ is a generator of $\widehat{\mathbb{F}_{p}^{\times}}$.
We now state our next lemma (cf. [3, Proposition 3.6.4]).
Lemma 4.2.
Let notations be as above. Then for any integers $1\leq i,j\leq p-2$,
$J_{p}(\omega_{\mathfrak{p}}^{-i},\omega_{\mathfrak{p}}^{-j})\equiv-\binom{i+j}{i}\ ({\rm{mod}}\ \mathfrak{p}).$
In particular, if $i+j\geq p$, then
$J_{p}(\omega_{\mathfrak{p}}^{-i},\omega_{\mathfrak{p}}^{-j})\equiv 0\ ({\rm{mod}}\ \mathfrak{p}).$
We also need the following known result (cf. [1, Theorem 2.1.4]).
Lemma 4.3.
Let $p>2$ be an odd prime. Then for any nontrivial character $\psi\in\widehat{\mathbb{F}_{p}^{\times}}$,
$J_{p}(\phi,\psi)=\psi(4)\cdot J_{p}(\psi,\psi).$
Now we are in a position to prove our last result.
@@
is a square of the integer $z_{p}$.
It remains to prove that the integer $z_{p}$ is a quadratic residue modulo $p$. Let notations be as in Lemma 4.2. As $(\frac{z_{p}}{p})$ is independent of the choice of the generator $\chi_{p}$, we now let $\chi_{p}=\omega_{\mathfrak{p}}$ be the Teichümuller character $\omega_{\mathfrak{p}}$ of $\mathfrak{p}$. Then by Lemma 4.3 and Lemma 4.2
$y_{p}$ $\equiv{\bm{i}}^{(n-2)/2}\cdot\sqrt{p-1}\cdot\prod_{k\in(0,n/2)}\left(J_{p}(\phi,\chi_{q}^{k})-J_{p}(\phi,\chi_{p}^{-k})\right)$
$\equiv{\bm{i}}^{(n-2)/2}\cdot\sqrt{p-1}\cdot\prod_{k\in(0,n/2)}-J_{p}(\phi,\chi_{p}^{-k})$
$\equiv{\bm{i}}^{(n-2)/2}\cdot\sqrt{p-1}\cdot\prod_{k\in(0,n/2)}-\chi_{p}^{-k}(4)\cdot J_{p}(\chi_{p}^{-k},\chi_{p}^{-k})$
$\equiv{\bm{i}}^{(n-2)/2}\cdot\sqrt{p-1}\cdot\prod_{k\in(0,n/2)}\chi_{p}^{-k}(4)\cdot\binom{2k}{k}\ ({\rm{mod}}\ \mathfrak{p}).$
By this and Lemma 4.1, one can verify that
$\left(\frac{z_{p}}{p}\right)=\left(\frac{2}{p}\right)^{\frac{n-2}{2}}\left(\frac{y_{p}}{p}\right)=\left(\frac{2}{p}\right)^{\frac{n}{2}}\left(\frac{\sqrt{-1}}{p}\right)^{\frac{n}{2}}=1.$
In view of the above, we have completed the proof. ∎
@@ 5. Concluding Remarks
for some intger $t_{p}$ with $(\frac{t_{p}}{p})=1$. On the other hand, by the proof of Corollary 1.1, we have proven that
@@
$z_{p}=\frac{{\bm{i}}^{\frac{p-5}{4}}\cdot I_{p}(\chi_{p})\cdot\sqrt{p-1}}{2^{(p-5)/4}}\in\mathbb{Z}$
with $(\frac{z_{p}}{p})=1$.
In view of the above, we clearly see that Conjecture 5.1 holds if and only if the rational number $4z_{p}/(p-1)$ is an integer. However, we can not prove this result by the methods used in this paper.